### v1

@@ 5.1.1 Accreting Millisecond X-ray Pulsars
The eXTP mission is suited to address these questions. For example, the high time resolution and sensitivity of the SFA-T will allow precise tracking of the pulse frequency evolution and phase-resolved spectroscopy to investigate the accretion geometry and the coupling of the disk-magnetosphere. In parallel, the PFA will provide direct measurements of X-ray polarization, giving access to emission geometry and magnetic field topology. The variation of polarization properties over the pulse phase will further constrain models of beam formation and magnetospheric scattering.
@@
Transitions between accretion and propeller states can be studied in detail by monitoring spectral and timing variability with the SFA. At the same time, changes in the angle or degree of polarization, measurable with the PFA, can reveal whether the emission geometry or the flow pattern has changed fundamentally [153]. Combined, these diagnostics provide a multi-dimensional view of state transitions.
@@
An intriguing recent findings in AMXP research is the possible presence of disk-driven outflows, observed via absorption features in their X-ray spectra [154]. These features, including blue-shifted Fe XXV and Fe XXVI lines near 7 keV, have traditionally been detected with long exposures of high-resolution spectrometers such as those on board XMM-Newton and Chandra. However, simulations using eXTP’s response matrices show that the SFA can also detect these lines in moderately deep exposures. In particular, simulated spectra for the eclipsing AMXP Swift J1749.4–2807, based on NICER observations and including a Gaussian absorption feature, demonstrate that with 15–25 ks exposures and six modules, eXTP can detect absorption lines a factor of three fainter than that observed in this source, and determine whether they are significantly blue-shifted. A representative fit is shown in Figure 7, where the contour plot and spectral model illustrate both detectability and constraints on blueshift. The PFA may also reveal complementary evidence for outflows by detecting scattering-induced polarization changes associated with extended structures, such as disk winds.
@@
From a spectral-timing perspective, the detection and characterization of iron K-shell transitions—especially Fe XXV and Fe XXVI near 6.7–7 keV, offer crucial diagnostics of plasma conditions near the neutron star. The SFA’s energy resolution (180 eV at 6 keV) and large effective area allow for the detection of such features even in short exposures (10–20 ks), enables phase-resolved spectroscopy to study how the line parameters vary across the pulsar spin phase. In sources with outflows, this can reveal changes in the ionization state and dynamics of the absorber on short timescales. Simultaneously, the high timing resolution (10 s) permits detailed modeling of pulse profiles and their energy dependence, allowing separation of Comptonized and thermal components and tracking of phase lags or frequency drifts. These joint capabilities will be especially powerful in bright AMXPs undergoing outbursts, where variations in the iron line strength or width may correlate with pulse amplitude, providing direct observational evidence of inner disk dynamics modulated by the star’s magnetic field.
@@ 5.1.2 Transitional Millisecond pulsars
Millisecond pulsars (MSPs, defined with spin period ms) are believed to be formed through the “recycling” scenario by accreting mass and angular momentum from a companion star in low-mass (donor mass ) X-ray binary (LMXB) systems in a Gyr-long phase [155, 156]. Currently, 552 MSPs have been detected in the Galactic field 111https://www.astro.umd.edu/~eferrara/pulsars/GalacticMSPs.txt, representing of the known Galactic pulsar population. In comparison, 80% of the 344 radio pulsars found in 45 globular clusters (GCs) are MSPs222https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html. This phenomenon has been attributed to the higher stellar density thus higher stellar encounter rate in GCs which resulted in more abundant LMXBs in GCs than in the field [157, 158, 159]. While the majority of MSPs are detected in radio with an increasing fraction in gamma rays333https://confluence.slac.stanford.edu/display/GLAMCOG/Public+List+of+LAT-Detected+Gamma-Ray+Pulsars, most of them being rotation powered [see the ATNF pulsar catalog444https://www.atnf.csiro.au/research/pulsar/psrcat/ 160], two dozen accreting MSPs (AMXPs) in X-ray transients have been found [161, 162]. The discovery of GC MSPs and AMXPs reasonably provided supports to the “recycling” scenario.
@@
Discovered just a decade ago, transitional MSPs (tMSPs) are defined by the switch between the accretion- and rotation-powered states back and forth over a few years. They are considered as the “missing link” between radio MSPs and LMXBs, providing conclusive proof that an evolutionary link exists between these two classes remain a hot topic on the frontier of astrophysics. So far, there are only three firmly identified tMSPs, PSR J1023+0038 [163], XSS J122704859 [164] and IGR J182452452 [165] and a few candidates recognized, representing a very limited sample. During the transitions, the X-ray luminosity changes by at least an order of magnitude, accompanied with multi-wavelengh emissions from radio, optical/UV to gamma rays [for a review, see 166]. The unique “two-sidedness” of tMSPs provides a natural laboratory for understanding neutron star evolution, accretion physics, and extreme magnetic field environments.
@@
Despite the progress made by a decade of observation and theoretical modeling of tMSPs, the underlying physics of tMSP is still not well understood and there are many open questions. With its exceptional large effective area, high time resolution, and precise polarization measurement capability, eXTP has unique advantages in studying the radiation characteristics and state changes of tMSPs, eventually advancing our knowledge of tMSPs by potentially solving the following key problems.
@@
What are the exact mechanisms that trigger the state transitions? In the rotation-powered state ( erg ), tMSPs behave as “redbacks”, a sub-catagory of MSPs in tight binaries ( day) with a companion mass . Redbacks and thought to “devour” their companions through ablation by the pulsar wind, producing synchrotron radiation which is observable at X-rays and gamma rays. Such sources exhibit significant radio eclipses due to the presence of intra-binary material. In the accretion-powered states, so far, the accretion outburst ( erg ) similar to other AMXPs has only been observed in IGR J182452452. The tMSPs preferentially show a peculiar intermediate state, termed “sub-luminous disc state” which is characterized by the presence of an accretion disc and erg [166, and references therein].
@@
eXTP is expected to perform high cadence and high statistic follow-up observations once the state transition is triggered by X-ray monitors, because of the unprecedented large effective area of eXTP-SFA. With high-resolution spectral and timing analysis, eXTP will be able to capture the abrupt and dynamic processes associated with the state transition, shedding light on the physical mechanisms that are onset during the transition.
@@
emission is variable and three intensity modes (high, low and flaring modes) have been identified from the X-ray light curves [167]. The tMSPs show unexpectedly fast mode transitions on a time scale of 10 s with modes duration ranging from a few tens of seconds to a few hours. In this state, X-ray, UV and optical pulsations from two of the tMSPs have been detected in the high mode [168, 169, 170]. Gamma-ray emission is several magnitudes brightened when switching from the rotation-powered state to the sub-luminous disc state, and vice versa, dimmer when turning into the rotation-powered state [166, and references therein].
@@
The current tMSP sample size is very limited, prohibiting a full understanding of the complex and mysterious phenomena observed. Discovering more tMSPs to enlarge the sample is not only very important to answer the previous two questions but also to explore whether the state transition can occur in all MSPs in tight binaries. So far, the three tMSPs are all redbacks which are mostly found through dedicated multi-wavelength observations of Fermi-LAT unidentified (UnID) sources [over 2200 UnID in the latest 4FGL-DR4 catalog, 181]. On the one hand, the eXTP can perform follow-up observations of current tMSPs whose state transitions are triggered by wide-field X-ray monitors.
However, the high eXTP sensitivity will allow targeting roughly twice as faint X-ray counterparts of UnID gamma-ray sources to look for the distinctive features of tMSPs in the “subluminous disc state” such as the high and low modes, to date the most efficient way to identify candidate tMSPs [e.g. 166, and references therein].
@@ 5.1.3 Weakly magnetized neutron star systems
Another important sub-class of LMXBs is that hosting a weakly magnetized neutron star (NS) with G. These systems have historically been divided into two further subclasses, the Z and atoll sources [182], on the basis of their temporal pattern in the color-color or hardness-intensity diagram (CCD and HID, respectively). In their brightest, high-accretion regimes (High Soft State, HSS), the spectra of LMXBs are usually described by a two-component model consisting of a strong dominating Comptonization spectrum at higher energies plus soft emission attributed either to the accretion disk [183, 184], or a direct blackbody (BB) at temperature keV coming from the NS [185]. The two scenarios were labeled Eastern Model (EM) and Western Model (WM), respectively. In many cases, in addition to the persistent two-component continuum, in both Z and atoll sources clear evidence of a reflection feature attributed to the accretion disk, which can be fully or partially ionized [186, 187, 188, 189]. All known Z sources (Cyg X-2, GX 5-1, GX 340+0, Sco X-1, GX 17+2 and GX 349+2) always remain the high soft state (HSS), where the Comptonization component has low temperature ( keV) and high optical depth 5–10 depending on the assumption of plasma geometry, slab-like or spherical. Atoll sources, on the other hand, can be divided into two further subclasses: sources of the first class (GX 13+1, GX 3+1, GX 9+9, GX 9+1) have only been observed in the HSS, while the others experience state transition from HSS to low hard state (LHS), characterized by a dimmer flux in the X-ray band, and a significantly harder spectrum where the electron temperature can achieve about 30 keV, in low optical depth environment [e.g., 190, 191, 192, 193].
@@
The decisive tool appears to be unequivocally polarimetry, as demonstrated since 2021 with the launch of the Imaging X-ray Polarimetry Explorer [IXPE, 195].
What IXPE has told us so far is that polarization increases with energy, which is only possible if the soft and hard components do not share the same polarization angle [196, 197, 198, 199, 200]. Then a boundary layer with high optical depth and coplanar to the accretion disk seems to be ruled out. However, because of IXPE’s statistics, the positive correlation of the PD with energy is visible qualitatively, but it is not possible to define a truly parameterizable trend for individual sources, and in most of the cases the results are presented for the 2-4 keV and 4-8 keV energy bands only.
@@ 5.1.4 Thermonuclear X-ray bursts
Thermonuclear X-ray bursts, also known as type I X-ray bursts [e.g. 203], occur as a result of thermonuclear runaways on the surface of a neutron star in a binary system. These bursts occur in the layers above the crust, where freshly accreted fuel, primarily composed of hydrogen and helium, can burn in an unsteady way when compressed to a density of around g cm-3. Unstable burning takes place when the cooling cannot balance the heating due to the nuclear reactions. Key factors that affect the burning and its stability are the accreted composition, accretion rate and the energy exchanged with the deeper layers, including the crust. Very importantly, the observed light curves are affected by relativistic effects, especially those due to the gravitational field of the star that determines the path and spectrum of the light emitted from the surface [e.g. 204].
Despite a general understanding of their origin, type I bursts present several unresolved questions, many of which are fundamental to astrophysics. Also, because of their nature and relation to the neutron star and its environment, the bursts can be used to gain insight into physics of particular relevance. First and foremost, because of the effects of the gravitational field of the star, which depends on its mass and radius, their light curves can be used to put constraints on the internal structure of neutron stars and the exotic nuclear processes occurring within. Other key areas of interest include the geometry of accretion flows, magnetic field interactions, and unusual stellar abundances [see e.g. 205]. Furthermore, the very phenomenology of the type I bursts still presents challenges to our theoretical modelling. For example, the critical mass accretion rate at which the bursts are quenched seems to be almost an order of magnitude smaller than predicted. This is probably related to the fluid dynamics on the surface, with spin and magnetic fields playing a major role, and to the nuclear reactions involved [e.g. 204, 206].
In order to extract as much information as possible from the short and bright burst light curves, we need high-precision observations such as those provided by advanced X-ray telescopes and eXTP is particularly well-suited for this task. As an example, in the left panel of Figure 10, we compare the simulated light curves of type I X-ray bursts as detected by NICER and SFA-T. With its exceptional sensitivity and timing resolution, eXTP will be able to capture the rapid and dynamic processes associated with thermonuclear bursts, shedding light on the underlying physics and helping to answer many of the open questions in this field.
@@
Burst oscillations are a fascinating phenomenon observed in approximately 10% of all type I X-ray bursts [207]. These oscillations occur at frequencies closely matching the neutron star’s spin frequency, ranging between 11 and 620 Hz. They appear as strong signals in the power spectra of the bursts, sometimes at constant frequency and sometimes slowly drifting by a few Hz during the burst evolution. They are thought to be due to asymmetries in the surface emission modulated by the rotation of the star and its gravitational field effects. Thus, they provide a valuable diagnostic tool for studying neutron star properties (since the pulse profiles are affected by its gravitational field) and the dynamics of the burst process (which determines the surface emission pattern) [for a review, see 208].
@@
The origin of burst oscillations during the cooling (or tail) phase remains a subject of debate. Theories that explain non-uniform emission include varying fuel depths across the neutron star surface and cooling wakes [e.g. 212]. An alternative and intriguing hypothesis involves the excitation of global surface modes triggered by instabilities in the burning ocean both during the propagation of the deflagration front and in the cooling phase. This idea was initially proposed by [213] and expanded upon by subsequent theoretical work [e.g. 214, 215, 216, 217].
eXTP will enable high-precision tracking of burst oscillation frequency drift, which is crucial for understanding the burning front propagation and oscillation mechanism. eXTP will also allow for the first high-statistics studies of time-dependent phase-resolved spectroscopy during bursts. By tracking the spectral evolution of the oscillation signal, we can distinguish between flame spreading, wave patterns, and other potential burst asymmetry mechanisms.
@@
According to theory, the burst rate depends mainly on the accretion rate. However, theory predicts a steady increase in the burst rate, while observationally it has been noted that many sources display a phase when the burst ignition rate decreases with the accretion rate. The reasons for this behavior are still unknown, but it has been suggested that it could be related to some varying conditions on the surface [219, 220, 206], or to additional sources of heat [221, 222, 223], including mysterious shallow heating [224]. Since the problem is related to the transition of the bursts to stable burning (see also below), another possibility is errors in the estimates for the reaction rates [e.g. 225, 226]. A combination of these effects should change the ignition latitude of the bursts, which could also have an effect on the shape of the rise [227] and the presence and shape of the burst oscillations. Timing studies with eXTP have the potential to resolve burst oscillations during the rapid rise phase, and thus provide the clues necessary to identifying the ignition latitude and the underlying mechanism behind this striking discrepancy between observations and theory. We performed simulations of type I X-ray bursts ignited at three latitudes on the surface of a neutron star spinning at 400 Hz. In the right panel of Figure 10 we show the resulting light curves of the burst rise phase, as well as the temporal evolution of the 400 Hz burst oscillation intensity, demonstrating how eXTP could distinguish the three cases.
@@
Marginally stable burning occurs in the transition between unstable and stable burning. In this regime, the heating rate is alternatively greater and smaller than the cooling rate, while remaining always very close [228], thus generating QPOs at a typical frequency of 5–10 mHz (mHz QPOs) in the soft X-ray band [e.g., 229, 228, 230]. The mHz QPOs occur within a narrow range of non-burst luminosities, keV (5–11) 1036 erg/s, and should be detected by eXTP with strong flux variations below 5 keV. The mHz QPOs usually, but not always, disappear right before a type I X-ray burst [e.g., 229, 231].
@@
Currently, the main question is how the mHz QPOs come about. It has been suggested that they are due to nonuniform, area changing, surface emission [232, 233] or to a global variation of the surface blackbody temperature [234]. Which scenario is the right one?
@@
The results supporting the former scenario are based on data from XMM-Newton, while the second interpretation is based on NICER data. The different conclusions probably stem from the fit parameter degeneracy between the blackbody temperature and its normalization, which is worst when the photon counts are low. Leveraging its large effective area and high observation efficiency, eXTP is anticipated to provide substantial photon statistics, thereby enabling us to put stringent constraints on the relevant parameters. For example, in Figure 11 we compare the results of a fit from simulated observations with SFA-T to those obtained by NICER. The 1- error bars are reduced by a factor of 2 to 3.
@@
Another key issue is how to trigger marginally stable burning at the observed low accretion rate (10%–50% Eddington) [e.g., 229, 231, 235] rather than the 100 % of Eddington rate predicted by simulations [228, 225]. Two scenarios have been proposed to bridge this accretion rate gap: (a) it is the local accretion rate of the burning region, instead of the global average rate of the entire neutron star, that triggers the marginally stable burning [228, 236]; (b) there is some extra heat flux from the crust and/or a turbulent mixing of helium in the deep layer, which could increase the stability of the burning and hence lower the critical rate for the stable burning [e.g. 222]. The validity of either scenario is still debated, and it is challenging to directly study the local accretion rate or turbulent mixing from observations. In the transitional spectral state, the frequency of the millihertz QPOs is observed to systematically decrease with time, until the QPOs disappear and an X-ray burst occurs. The turbulent mixing scenario predicts an increase in the oscillation amplitude along with this drift. Thus, precisely measuring the evolution of the amplitude is a possible way to test this scenario. With its large effective area in the soft X-ray band, eXTP is promising in identifying these possible subtle amplitude variations and hence establishing the validity of turbulent mixing theory.
@@
A further conundrum is the coexistence of mHz QPOs and type I bursts in some sources [e.g., 229, 231, 237], since the two phenomena should appear sequentially [228]. A possible explanation can again be the different local conditions on the surface of the star [220, 236, 206]. The spectral-timing capabilities of eXTP can allow us to study in greater detail both the persistent and the burst/mHz QPOs emission properties, giving clues to identify which solution best explains the observations.

### v2

@@ 5.1.1 Accreting Millisecond X-ray Pulsars
The eXTP mission is suited to address these questions. For example, the high time resolution and large effective area of the SFA-T will allow precise tracking of the pulse frequency evolution and phase-resolved spectroscopy to investigate the accretion geometry and the coupling of the disk-magnetosphere. In parallel, the PFA will provide direct measurements of X-ray polarization, giving access to emission geometry and magnetic field topology. The variation of polarization properties over the pulse phase will further constrain models of beam formation and magnetospheric scattering.
@@
Transitions between accretion and propeller states can be studied in detail by monitoring spectral and timing variability with the SFA. At the same time, changes in the angle or degree of polarization, measurable with the PFA, can reveal whether the emission geometry or the flow pattern have changed fundamentally [154]. Combined, these diagnostics provide a multi-dimensional view of state transitions.
@@
An intriguing recent findings in AMXP research is the possible presence of disk-driven outflows, observed via absorption features in their X-ray spectra [155]. These features, including blue-shifted Fe XXV and Fe XXVI lines near 7 keV, have traditionally been detected with long exposures of high-resolution spectrometers such as those on board XMM-Newton and Chandra. However, simulations using eXTP’s response matrices show that the SFA can also detect these lines in moderately deep exposures. In particular, simulated spectra for the eclipsing AMXP Swift J1749.4–2807, based on NICER observations and including a Gaussian absorption feature, demonstrate that with 15–25 ks exposures and six SFA modules, eXTP can detect absorption lines a factor of three fainter than that observed in this source, and determine whether they are significantly blue-shifted. A representative fit is shown in Figure 7, where the contour plot and spectral model illustrate both detectability and constraints on blueshift. The PFA may also reveal complementary evidence for outflows by detecting scattering-induced polarization changes associated with extended structures, such as disk winds.
@@
From a spectral-timing perspective, the detection and characterization of iron K-shell transitions—especially Fe XXV and Fe XXVI near 6.7–7 keV, offer crucial diagnostics of plasma conditions near the neutron star. The SFA’s energy resolution (180 eV at 6 keV) and large effective area allow for the detection of such features even in short exposures (10–20 ks), enable phase-resolved spectroscopy to study how the line parameters vary across the pulsar spin phase. In sources with outflows, this can reveal changes in the ionization state and dynamics of the absorber on short timescales. Simultaneously, the high timing resolution (10 s) permits detailed modeling of pulse profiles and their energy dependence, allowing separation of Comptonized and thermal components and tracking of phase lags or frequency drifts. These joint capabilities will be especially powerful in bright AMXPs undergoing outbursts, where variations in the iron line strength or width may correlate with pulse amplitude, providing direct observational evidence of inner disk dynamics modulated by the star’s magnetic field.
@@ 5.1.2 Transitional Millisecond pulsars
Millisecond pulsars (MSPs, defined with spin period ms) are believed to be formed through the “recycling” scenario by accreting mass and angular momentum from a companion star in LMXB systems in a Gyr-long phase [156, 157]. Currently, 552 MSPs have been detected in the Galactic field 111https://www.astro.umd.edu/~eferrara/pulsars/GalacticMSPs.txt, representing of the known Galactic pulsar population. In comparison, 80% of the 344 radio pulsars found in 45 globular clusters (GCs) are MSPs222https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html. This phenomenon has been attributed to the higher stellar density thus higher stellar encounter rate in GCs which resulted in more abundant LMXBs in GCs than in the field [158, 159, 160]. While the majority of MSPs are detected in radio with an increasing fraction in gamma rays333https://confluence.slac.stanford.edu/display/GLAMCOG/Public+List+of+LAT-Detected+Gamma-Ray+Pulsars, most of them being rotation powered [see the ATNF pulsar catalog444https://www.atnf.csiro.au/research/pulsar/psrcat/, 161], two dozen accreting MSPs (AMXPs) in X-ray transients have been found [162, 163]. The discovery of GC MSPs and AMXPs reasonably provided supports for the “recycling” scenario.
@@
Discovered just a decade ago, transitional MSPs (tMSPs) are defined by the switch between the accretion- and rotation-powered states back and forth over a few years. They are considered as the “missing link” between radio MSPs and LMXBs, providing conclusive proof of an evolutionary link between these two classes. So far, there are only three firmly identified tMSPs, PSR J1023+0038 [164], XSS J122704859 [165] and IGR J182452452 [166] and a few candidates recognized, representing a very limited sample. During the transitions, the X-ray luminosity changes by at least an order of magnitude, accompanied with multi-wavelengh emissions from radio, optical/UV to gamma rays [for a review, see 167]. The unique “two-sidedness” of tMSPs provides a natural laboratory for understanding neutron star evolution, accretion physics, and extreme magnetic field environments.
@@
Despite the progress made by a decade of observation and theoretical modeling of tMSPs, the underlying physics of tMSP is still not well understood and there are many open questions. With its exceptional large effective area, high time resolution, and polarization measurement capability, eXTP has unique advantages in studying the radiation characteristics and state changes of tMSPs, eventually advancing our knowledge of tMSPs by potentially solving the following key problems.
@@
What are the exact mechanisms that trigger the state transitions? In the rotation-powered state ( erg ), tMSPs behave as “redbacks”, a sub-catagory of MSPs in tight binaries ( day) with a companion mass . Redbacks are thought to “devour” their companions through ablation by the pulsar wind, producing synchrotron radiation which is observable at X-rays and gamma rays. Such sources exhibit significant radio eclipses due to the presence of intra-binary material. In the accretion-powered states, so far, the accretion outburst ( erg ) similar to other AMXPs has only been observed in IGR J182452452. The tMSPs preferentially show a peculiar intermediate state, termed “sub-luminous disc state” which is characterized by the presence of an accretion disc and erg [167, and references therein].
@@
eXTP is expected to perform high cadence and high statistic follow-up observations once the state transition is triggered by X-ray monitors, because of the unprecedented large effective area of eXTP-SFA. With spectral and high-resolution timing analysis, eXTP will be able to capture the abrupt and dynamic processes associated with the state transition, shedding light on the physical mechanisms that are onset during the transition.
@@
emission is variable and three intensity modes (high, low and flaring modes) have been identified from the X-ray light curves [168]. The tMSPs show unexpectedly fast mode transitions on a time scale of 10 s with modes duration ranging from a few tens of seconds to a few hours. In this state, X-ray, UV and optical pulsations from two of the tMSPs have been detected in the high mode [169, 170, 171]. Gamma-ray emission is several magnitudes brightened when switching from the rotation-powered state to the sub-luminous disc state, and vice versa, dimmer when turning into the rotation-powered state [167, and references therein].
@@
The current tMSP sample size is very limited, prohibiting a full understanding of the complex and mysterious phenomena observed. Discovering more tMSPs to enlarge the sample is not only very important to answer the previous two questions but also to explore whether the state transition can occur in all MSPs in tight binaries. So far, the three tMSPs are all redbacks which are mostly found through dedicated multi-wavelength observations of Fermi-LAT unidentified (UnID) sources [over 2200 UnID in the latest 4FGL-DR4 catalog, 182]. On the one hand, the eXTP can perform follow-up observations of current tMSPs whose state transitions are triggered by wide-field X-ray monitors.
However, the high eXTP sensitivity will allow targeting roughly twice as faint X-ray counterparts of UnID gamma-ray sources to look for the distinctive features of tMSPs in the “subluminous disc state” such as the high and low modes, to date the most efficient way to identify candidate tMSPs [e.g. 167, and references therein].
@@ 5.1.3 Weakly magnetized neutron star systems
Another important sub-class of LMXBs is that hosting a weakly magnetized neutron star (NS) with G. These systems have historically been divided into two further subclasses, the Z and atoll sources [183], on the basis of their temporal pattern in the color-color or hardness-intensity diagram (CCD and HID, respectively). In their brightest, high-accretion regimes (High Soft State, HSS), the spectra of LMXBs are usually described by a two-component model consisting of a strong dominating Comptonization spectrum at higher energies plus soft emission attributed either to the accretion disk [184, 185], or a direct blackbody (BB) at temperature keV coming from the NS [186]. The two scenarios were labeled Eastern Model (EM) and Western Model (WM), respectively. In many cases, in addition to the persistent two-component continuum, in both Z and atoll sources clear evidence of a reflection feature attributed to the accretion disk, which can be fully or partially ionized [187, 188, 189, 190]. All known Z sources (Cyg X-2, GX 5-1, GX 340+0, Sco X-1, GX 17+2 and GX 349+2) always remain the high soft state (HSS), where the Comptonization component has low temperature ( keV) and high optical depth 5–10 depending on the assumption of plasma geometry, slab-like or spherical. Atoll sources, on the other hand, can be divided into two further subclasses: sources of the first class (GX 13+1, GX 3+1, GX 9+9, GX 9+1) have only been observed in the HSS, while the others experience state transition from HSS to low hard state (LHS), characterized by a dimmer flux in the X-ray band, and a significantly harder spectrum where the electron temperature can achieve about 30 keV, in low optical depth environment [e.g., 191, 192, 193, 194].
@@
The decisive tool appears to be unequivocally polarimetry, as demonstrated since 2021 with the launch of the IXPE [196]. What IXPE has told us so far is that polarization increases with energy, which is only possible if the soft and hard components do not share the same polarization angle [197, 198, 199, 200, 201]. Then a boundary layer with high optical depth and coplanar to the accretion disk seems to be ruled out. However, because of IXPE’s statistics, the positive correlation of the PD with energy is visible qualitatively, but it is not possible to define a truly parameterizable trend for individual sources, and in most of the cases the results are presented for the 2-4 keV and 4-8 keV energy bands only.
@@ 5.1.4 Thermonuclear X-ray bursts
Thermonuclear X-ray bursts, also known as type I X-ray bursts [e.g. 204], occur as a result of thermonuclear runaways on the surface of a neutron star in a binary system. These bursts occur in the layers above the crust, where freshly accreted fuel, primarily composed of hydrogen and helium, can burn in an unsteady way when compressed to a density of around g cm-3. Unstable burning takes place when the cooling cannot balance the heating due to the nuclear reactions. Key factors that affect the burning and its stability are the accreted composition, accretion rate and the energy exchanged with the deeper layers, including the crust. Very importantly, the observed light curves are affected by relativistic effects, especially those due to the gravitational field of the star that determines the path and spectrum of the light emitted from the surface [e.g. 205].
Despite a general understanding of their origin, type I bursts present several unresolved questions, many of which are fundamental to astrophysics. Also, because of their nature and relation to the neutron star and its environment, the bursts can be used to gain insight into physics of particular relevance. First and foremost, because of the effects of the gravitational field of the star, which depends on its mass and radius, their light curves can be used to put constraints on the internal structure of neutron stars and the exotic nuclear processes occurring within. Other key areas of interest include the geometry of accretion flows, magnetic field interactions, and unusual stellar abundances [see e.g. 206]. Furthermore, the very phenomenology of the type I bursts still presents challenges to our theoretical modelling. For example, the critical mass accretion rate at which the bursts are quenched seems to be almost an order of magnitude smaller than predicted. This is probably related to the fluid dynamics on the surface, with spin and magnetic fields playing a major role, and to the nuclear reactions involved [e.g. 205, 207].
In order to extract as much information as possible from the short and bright burst light curves, we need high-precision observations such as those provided by advanced X-ray telescopes and eXTP is particularly well-suited for this task. As an example, in the left panel of Figure 10, we compare the simulated light curves of type I X-ray bursts as detected by NICER and SFA. With its exceptional sensitivity and timing resolution, eXTP will be able to capture the rapid and dynamic processes associated with thermonuclear bursts, shedding light on the underlying physics and helping to answer many of the open questions in this field.
@@
Burst oscillations are a fascinating phenomenon observed in approximately 10% of all type I X-ray bursts [208]. These oscillations occur at frequencies closely matching the neutron star’s spin frequency, ranging between 11 and 620 Hz. They appear as strong signals in the power spectra of the bursts, sometimes at constant frequency and sometimes slowly drifting by a few Hz during the burst evolution. They are thought to be due to asymmetries in the surface emission modulated by the rotation of the star and its gravitational field effects. Thus, they provide a valuable diagnostic tool for studying neutron star properties (since the pulse profiles are affected by its gravitational field) and the dynamics of the burst process (which determines the surface emission pattern) [for a review, see 209].
@@
The origin of burst oscillations during the cooling (or tail) phase remains a subject of debate. Theories that explain non-uniform emission include varying fuel depths across the neutron star surface and cooling wakes [e.g. 213]. An alternative and intriguing hypothesis involves the excitation of global surface modes triggered by instabilities in the burning ocean both during the propagation of the deflagration front and in the cooling phase. This idea was initially proposed by [214] and expanded upon by subsequent theoretical work [e.g. 215, 216, 217, 218].
eXTP-SFA will enable high-precision tracking of burst oscillation frequency drift, which is crucial for understanding the burning front propagation and oscillation mechanism. eXTP will also allow for the first high-statistics studies of time-dependent phase-resolved spectroscopy during bursts. By tracking the spectral evolution of the oscillation signal, we can distinguish between flame spreading, wave patterns, and other potential burst asymmetry mechanisms.
@@
According to theory, the burst rate depends mainly on the accretion rate. However, theory predicts a steady increase in the burst rate, while observationally it has been noted that many sources display a phase when the burst ignition rate decreases with the accretion rate. The reasons for this behavior are still unknown, but it has been suggested that it could be related to some varying conditions on the surface [220, 221, 207], or to additional sources of heat [222, 223, 224], including mysterious shallow heating [225]. Since the problem is related to the transition of the bursts to stable burning (see also below), another possibility is errors in the estimates for the reaction rates [e.g. 226, 227]. A combination of these effects should change the ignition latitude of the bursts, which could also have an effect on the shape of the rise [228] and the presence and shape of the burst oscillations. Timing studies with eXTP have the potential to resolve burst oscillations during the rapid rise phase, and thus provide the clues necessary to identifying the ignition latitude and the underlying mechanism behind this striking discrepancy between observations and theory. We performed simulations of type I X-ray bursts ignited at three latitudes on the surface of a neutron star spinning at 400 Hz. In the right panel of Figure 10 we show the resulting light curves of the burst rise phase, as well as the temporal evolution of the 400 Hz burst oscillation intensity, demonstrating how eXTP-SFA could distinguish the three cases.
@@
Marginally stable burning occurs in the transition between unstable and stable burning. In this regime, the heating rate is alternatively greater and smaller than the cooling rate, while remaining always very close [229], thus generating QPOs at a typical frequency of 5–10 mHz (mHz QPOs) in the soft X-ray band [e.g., 230, 229, 231]. The mHz QPOs occur within a narrow range of non-burst luminosities, (5–11) 1036 erg s-1, and should be detected by the SFA with strong flux variations below 5 keV. The mHz QPOs usually, but not always, disappear right before a type I X-ray burst [e.g., 230, 232].
@@
Currently, a crucial question is how the mHz QPOs come about. It has been suggested that they are due to nonuniform, area changing, surface emission [233, 234] or to a global variation of the surface blackbody temperature [235]. Which scenario is correct?
@@
The results supporting the former scenario are based on data from XMM-Newton, while the second interpretation is based on NICER data. The different conclusions probably stem from the fit parameter degeneracy between the blackbody temperature and its normalization. The degeneracy is strong when the photon counts are low. Leveraging its large effective area, eXTP-SFA is anticipated to provide substantial photon statistics, thereby enabling us to put stringent constraints on the relevant parameters. For example, Figure 11 compares the fit results from simulated observations with 5 SFA-T modules to those obtained by NICER. The 1- error bars are reduced by a factor of 2 to 3.
@@
Another key issue is how to trigger marginally stable burning at the observed low accretion rate (10%–50% Eddington) [e.g., 230, 232, 236] rather than the 100 % of Eddington rate predicted by simulations [229, 226]. Two scenarios have been proposed to bridge this accretion rate gap: (a) it is the local accretion rate of the burning region, instead of the global average rate of the entire neutron star, that triggers the marginally stable burning [229, 237]; (b) there is some extra heat flux from the crust and/or a turbulent mixing of helium in the deep layer, which could increase the stability of the burning and hence lower the critical rate for the stable burning [e.g. 223]. The validity of either scenario is still debated, and it is challenging to directly study the local accretion rate or turbulent mixing from observations. In the transitional spectral state, the frequency of the millihertz QPOs is observed to systematically decrease with time, until the QPOs disappear and an X-ray burst occurs. The turbulent mixing scenario predicts an increase in the oscillation amplitude along with this drift. Thus, precisely measuring the evolution of the amplitude is a possible way to test this scenario. With its large effective area in the soft X-ray band, the SFA is promising in identifying these possible subtle amplitude variations and hence establishing the validity of turbulent mixing theory.
@@
A further conundrum is the coexistence of mHz QPOs and type I bursts in some sources [e.g., 230, 232, 238], since the two phenomena should appear sequentially [229]. A possible explanation can again be the different local conditions on the surface of the star [221, 237, 207]. The spectral-timing capabilities of eXTP-SFA can allow us to study in greater detail both the persistent and the burst/mHz QPOs emission properties, giving clues to identify which solution best explains the observations.